A student is holding a stack of books with a combined mass of 20 kg while walking 5 m from her desk to the door. How much work is done (g=9.8 m/s^2)? Hint: What is the direction of the force and what is the direction of the displacement?
Answer #1

W = F*d*cos\theta

W= m*g*d*cos(0)

W= 20*9.8*5*1

W = 980 N